OpenStudy (anonymous):

Combining like terms... 12r-8-12. = -8r is this correct

OpenStudy (anonymous):

When combining like terms you just combine numbers with the same variables. For example: 2r+3r-7+1 If I combine like terms I get 2r+3r=5r and -7+1=-6 So you get 5r-6

OpenStudy (anonymous):

So since you have 12r-8-12 You have just 12r then -8-12=-20 So you get 12r-20

OpenStudy (anonymous):

Understand it now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks a lot!